OpenStudy (anonymous):

On a circle of radius 10 feet, what angle would subtend an arc of length 1 feet?

OpenStudy (anonymous):

We have this formula : \[l=R\theta \] Where : \[l~:~\text{the length of the arc }\\\theta~:~\text{the angle in radian }\\R~:~\text{the radius } \] So : \[\theta=\frac{l}R=\frac{1}{10}=0.1\text{rd}\] And you can convert it to degrees if you want !

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

I do not remember learning this

OpenStudy (anonymous):

You know that : \[180^\circ =\pi\text{rd}\] So : \[1\text{rd}=\frac{1}{\pi}180^\circ\] So : \[0.1\text{ rd}=0.1\times180^\circ\div\pi\approx 5.73^\circ\]

OpenStudy (anonymous):

what is r and what is d?

OpenStudy (anonymous):

rd means radian

OpenStudy (anonymous):

oh